OpenStudy (devonhoward15):

Find the roots of the polynomial question. will fan and medal x^2+24=11x

OpenStudy (devonhoward15):

@FaiqRaees

OpenStudy (faiqraees):

Start by giving it a quadratic form

OpenStudy (devonhoward15):

ok

OpenStudy (devonhoward15):

@Bobo-i-bo

OpenStudy (devonhoward15):

i got 3,8

OpenStudy (devonhoward15):

@FaiqRaees

OpenStudy (devonhoward15):

@welshfella

OpenStudy (faiqraees):

correct

OpenStudy (devonhoward15):

thanks

MsBrains (ms-brains):

I think you are in the wrong section here buddy. xP You should be posting this in math, not earth sciences.

OpenStudy (zwergje):

Roots: y= o x^2+24=11x x^2-11x+24 = 0 Use ABC formula to find x x = (-b + OR - root(b square - 4ac))/2a x = (11 + OR - root(-11 square - 4*1*24))/2*1 x = (11 + OR - root(121 - 96))/2 x = (11 + OR - root 25)/2 x = (11 + OR - 5)/2 x1 = (11 + 5)/2 = 8 x2 = (11 - 5)/2 = 3